OpenStudy (vishweshshrimali5):

Using LMVT (Lagrange's Mean Value Theorem) for solving algebra problems

OpenStudy (abb0t):

@dan815

OpenStudy (vishweshshrimali5):

The square roots of two successive natural numbers greater than \(\large{N^2}\) differ by: --------------------------------------------------------------------- Let n be a natural number greater than \(\large{N^2}\). Consider \(\large{f(x) = \sqrt{x}}\) where \(\large{x \in [n,n+1]}\). Since, f(x) satisfies conditions for LMVT, thus: \[\large{f'(c) = \cfrac{f(n+1)-f(n)}{n+1-n}}\] for some \(\large{c \in (n, n+1)}\tag{1}\) and \(\large{f'(x) = \cfrac{1}{2\sqrt{x}}}\) \(\large{\implies f'(c) = \cfrac{1}{2\sqrt{c}}}\tag{2}\) Thus, using (1) and (2): \[\large{\cfrac{1}{2\sqrt{c}} = \sqrt{n+1}-\sqrt{n}}\tag{3}\] Since, \(\large{n > N^2}\), thus: \[\large{c > N^2 ~\implies~\cfrac{1}{c} < \cfrac{1}{N^2} ~\implies~\cfrac{1}{2\sqrt{c}} < \cfrac{1}{2N}}\] Thus: \[\large{\boxed{\sqrt{n+1}-\sqrt{n} < \cfrac{1}{2N}}}\tag{Ans}\]

OpenStudy (vishweshshrimali5):

@ganeshie8 @Kainui @Mimi_x3 @satellite73 Any more ideas ?

OpenStudy (vishweshshrimali5):

This was a question of sequences and series and I used LMVT :P ;)

OpenStudy (kainui):

Interesting, I'd like to see what kind of applications this could have. I've actually never heard of Lagrange's mean value theorem before but it looks familiar.

OpenStudy (vishweshshrimali5):

It is a generalization of `Rolle's theorem`

OpenStudy (vishweshshrimali5):

http://en.wikipedia.org/wiki/Mean_value_theorem

OpenStudy (vishweshshrimali5):

@ikram002p Any idea ???